OpenStudy (anonymous):

Convert z = (- 5 √3 )/ 2 + (5/2)i to polar coordinates

jagr2713 (jagr2713):

is there an answer choice

OpenStudy (anonymous):

No. I have to show work and idk how to do these

OpenStudy (anonymous):

let z=\[\left( \frac{ -5\sqrt{3} }{ 2}+\iota \frac{ 5 }{ 2 } \right)=\left( r \cos \theta +\iota~ r \sin \theta \right)\] \[r~\cos \theta=\frac{ -5\sqrt{3} }{ 2 }....(1)\] \[r~\sin \theta=\frac{ 5 }{ 2 } ...(2)\] square and add \[r^2 \left( \cos ^2 \theta +\sin ^2 \theta \right)=\frac{ 75 }{ 4 }+\frac{ 25 }{ 4 }=\frac{ 75+25 }{ 4 }=\frac{ 100 }{ 4 }=25\] r=5 \[5 ~\cos \theta=\frac{ -5\sqrt{3} }{ 2 },\cos \theta=-\frac{ \sqrt{3} }{ 2 }=-\cos \frac{ \pi }{ 6 }=\cos \left( \pi-\frac{ \pi }{ 6 } \right)=\cos \frac{ 5~\pi } { 6 }\] \[\theta=\frac{ 5~\pi }{ 6 }\] \[z=\left( 5,\frac{ 5~\pi }{ 6 } \right)\]